-3x+y=-3 4x+2y=14 solved in substitution



Found 2 solutions by Alan3354, addingup:
Answer by Alan3354(69443)

-3x+y=-3 --> y = 3x-3
4x+2y=14 --> 2x+y = 7
----------
2x+y = 7
Sub for y
2x +(3x-3) = 7
5x = 10
x = 2
---
y = 3

Answer by addingup(3677)

-3x+y=-3; y = 3x-3
4x+2y=14; 2y = -4x+14; y = -2x+7
:
Now I've got these two equations in slope intercept form. Is there anything else they are asking you to do?